Circular No. 6530

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 C. Kouveliotou, Universities Space Research Association (USRA) and

Marshall Space Flight Center (MSFC), NASA; K. J. Deal, G. A. Richardson, and M. S. Briggs, University of Alabama in Huntsville (UAH); G. J. Fishman, MSFC; J. van Paradijs, University of Amsterdam and UAH, report: “We have recorded a series of outbursts starting on 1996 Dec. 2 that are consistent with the position of GRO J1744–28, the bursting pulsar. The bursts are detected between 20 and 100 keV, and on Dec. 2 they came in variable rates ranging from every 300 to every 500 s, with a preliminary total number estimated between 80 and 100 events. After Dec. 3, the burst rate dropped to 15–20 bursts/day, where it has remained. The current average error radius of the burst locations is ∼ 4^{o}. The burst fluence is currently ∼ 2 × 10^{−7} erg cm^{−2}. This indicates that the source is in outburst again. Although the source location is currently very close to the sun, we encourage observations in other wavelengths whenever possible.”

M. H. Finger and C. R. Robinson, USRA; B. A. Harmon, MSFC; and B. A. Vaughan, California Institute of Technology, report for the Compton Observatory BATSE team: “Persistent, hard x-ray emission from GRO J1744–28 (IAUC 6272, 6284, 6285) is currently being detected through both pulsation and earth occultation. Pulsations with a period near 0.467 s became detectable on Dec. 17, rising to a 20- to 40-keV-rms-pulsed flux of 1.3(1) × 10^{−9} erg cm^{−1} s^{−1} (170 mCrab) by 1997 Jan. 3. On Jan. 1.0, the intrinsic pulse frequency was 2.1408983(4) Hz, and the spin-up rate was 2.8(5) × 10^{−12} Hz s^{−1}. This assumes the following circular orbit parameters: P = 11.83665(14) days; epoch of longitude 90^{o} = JD 2450126.9977(4) TDB; a sin i = 2.6371(5) light-s (determined using data from the 1995 Dec.–1996 Apr. outburst). Earth-occultation monitoring of flux from GRO J1744–28 shows its intensity increasing by a factor of about two during 1996 Dec. 25–1997 Jan. 6. Although source confusion prevents a precise flux estimate, subtracting a 25-percent background for unresolved galactic-center sources yields a persistent flux for Jan. 3–6 in the energy bands 20–30, 30–40, and 40–50 keV of 590, 470, and 290 mCrab, respectively.”

COMET C/1995 O1 (HALE-BOPP) Visual m_{1} estimates with 10×50 binoculars: 1996 Dec. 14.70 UT, 4.2

(G. Glitscher, Darmstadt, Germany); 20.94, 4.2 (J. E. Bortle, Stormville, NY); 25.71, 3.8 (R. J. Bouma, Groningen, The Netherlands); 1997 Jan. 1.70, 3.3 (B. H. Granslo, near Oslo, Norway). 6.72, 3.1 (Granslo).

1997 January 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green